how many positive integers less than or equal to 70,are relatively prime 70?
Answer by nabla(475)

Use Euler's phi function. IE phi(n)=n(1-p_1)(1-p_2)...(1-p_r) where each p_i is a constituent prime (powers do not matter).
70=35*2=7*5*2
phi(70)=70(6/7)(4/5)(1/2)=6*4=24.